# From chimeras to extensive chaos in networks of heterogeneous Kuramoto oscillator populations

### v1

@@ Appendix D Antiphase states in ring networks
Here we provide the stablity analysis of antiphase states in topologies provided by either Eq.(16) or Eq. (17) with an even number of populations $N$. This analysis is performed in close analogy to that performed in Section III. Antiphase states are solutions of Eq. (3) given by
@@
Inserting this expressions into Eq. (3) and using the corresponding expression of $c_{\sigma\kappa}$ we obtain that
@@
Analogously to the stability of the homogeneous states in Section III, we apply the MSF formalism based on decomposing a perturbation on the basis of the eigenvectors of the connectivity matrix $C$. As a result we obtain that the stability of the antiphase state is provided by the eigenvalues of the $N$ $2\times 2$ matrices:
@@
where $\Lambda_{k}$ are provided by Eq. (18). The explicit expression of the system eigenvalues read:
@@
  • For $N>2$ other eigenvalues rather than $\Lambda_{1}$ can lead to instabilities of the antiphase state. This is the case of Fig. 4, in which the left branch of the instability is always given by $\Lambda_{1}$, but the right branch corresponds to other modes depending on $N$.
@@ Appendix E Transverse instabilities of the incoherent state
Here study transverse instabilities of the incoherent homogeneous state $Z_{\sigma}=0=:Z^{(0)}$ for $\sigma=1,\dots,N$. Our aim is to show that transverse instabilities do not arise when this state is stable, i.e., when $0<K\cos(\alpha)<2\Delta$. Without loss of generality, we impose $\overline{\omega}=0$. Then, by linearizing Eq. (3) around the incoherent solution $Z^{(0)}$ we obtain
$\dot{z}_{\sigma}$ $=-\Delta\omega z_{\sigma}+\frac{K}{2}\sum_{\kappa=1}^{M}c_{\sigma\kappa}z_{\kappa}e^{-i\alpha}=A^{(0)}z_{\sigma}+B^{(0)}\sum_{\kappa=1}^{N}c_{\sigma\kappa}z_{\kappa}\;,$ (37)
@@
The incoherent solution $Z^{(0)}$ is stable within the homogeneous manifold Eq. (4) if and only if $K\cos(\alpha)<2\Delta$. In this case, since $00<\Lambda_{k}<$ 0 and $\Delta>0$, we have that $\text{Tr}(M^{(0)}_{k})<0$. Therefore, transverse instabilities of the incoherent state may emerge only if it is already unstable to homogeneous perturbations.

### v2

@@ Appendix D Antiphase states in ring networks
Here we provide the stablity analysis of antiphase states in topologies given by either Eq.(16) or Eq. (17) with an even number of populations $N$. This analysis is performed in close analogy to that performed in Section III. Antiphase states are solutions of Eq. (3) given by
@@
Inserting these expressions into Eq. (3) and using the corresponding expression of $c_{\sigma\kappa}$ we obtain that
@@
Analogously to the stability of the homogeneous states shown in Section III, we apply the MSF formalism based on decomposing a perturbation on the basis of the eigenvectors of the connectivity matrix $C$. As a result we obtain that the stability of the antiphase state is given by the eigenvalues of the $N$ $2\times 2$ matrices:
@@
where $\Lambda_{k}$ are obtained from Eq. (18). The explicit expression of the system eigenvalues read:
@@
  • For $N>2$ other eigenvalues rather than $\Lambda_{1}$ can lead to instabilities of the antiphase state. This is the case of Fig. 4(a), in which the left branch of the instability is always given by $\Lambda_{1}$, but the right branch corresponds to other modes depending on $N$.
@@ Appendix E Transverse instabilities of the incoherent state
Here study the transverse instabilities of the incoherent homogeneous state $Z_{\sigma}=0$ for $\sigma=1,\dots,N$. Our aim is to show that transverse instabilities do not arise when this state is stable, i.e., when $0<K\cos(\alpha)<2\Delta$. Without loss of generality, we impose $\overline{\omega}=0$. Then, by linearizing Eq. (3) around the incoherent solution ($Z_{\sigma}=0$) we obtain
$\dot{z}_{\sigma}$ $=-\Delta\omega z_{\sigma}+\frac{K}{2}\sum_{\kappa=1}^{M}c_{\sigma\kappa}z_{\kappa}e^{-i\alpha}$ (37)
$=A^{(0)}z_{\sigma}+B^{(0)}\sum_{\kappa=1}^{N}c_{\sigma\kappa}z_{\kappa}\;,$
@@
The incoherent solution $Z^{(0)}$ is stable within the homogeneous manifold Eq. (4) if and only if $K\cos(\alpha)<2\Delta$. In this case, since $-1<\Lambda_{k}<1$ and $\Delta>0$, we have that $\text{Tr}(M^{(0)}_{k})<0$. Therefore, transverse instabilities of the incoherent state may emerge only if this state is already unstable to the homogeneous perturbation.